OpenStudy (anonymous):

let f(x)=(6x-7)^5 and find f"(x)

OpenStudy (anonymous):

have you differentiate it yet?

OpenStudy (anonymous):

\[f(x) = (6x-7)^5\] \[f'(x) = 5(6x-7)^4*6\]

OpenStudy (anonymous):

\[ => f'(x)= 30*(6x-7)^4\]

OpenStudy (anonymous):

yeh u did the first step

OpenStudy (anonymous):

now its your turn to differentiate it again :)

OpenStudy (anonymous):

use the product rule

OpenStudy (anonymous):

could we say diff like this 120(6x-7)^3

OpenStudy (anonymous):

is there something left for us knw

OpenStudy (anonymous):

\[f'(x) = 30*(6x-7)^4\] \[ f"(x) = 120(6x-7)^3*6\]

OpenStudy (anonymous):

i suppose you can do it like that

OpenStudy (anonymous):

that was suppose to be \(f''(x)=....\)

OpenStudy (anonymous):

i argued with a friends , they said it will stop up to 720(6x-7)^2

OpenStudy (anonymous):

\[f"(x) = 720 (6x-7)^3\]

OpenStudy (anonymous):

am i right up to this point Omni

OpenStudy (anonymous):

wolfram agrees with me; its not 720(6x-7)^2

OpenStudy (anonymous):

do u mean it will stop here 720(6x-7)^2

OpenStudy (anonymous):

that is not right

OpenStudy (anonymous):

the answer i wrote is right and wolfram agrees with me; there is a low chance that me and wolfram are wrong

OpenStudy (anonymous):

no no i just believe in u and am thankfull to u and ur is right